A 1200kg car is travelling at 20m/s when the driver slams on the brakes.
Determine the car's stopping distance if the static friction is 0.60?
Determine the car's stopping distance if the kinetic friction is 0.40?
1
Expert's answer
2020-04-06T09:04:05-0400

As the car is in motion then here static friction will act on it and it will provide a de-acceleration till the car does not stop

hence the deacceleration caused by the static friction is equal to "\mu g" = 4 m/s2


Applying formula here "v^2-u^2=2aS"

final velocity being zero and initial velocity is 20 m/s and a=-4 m/s2

"0-20^2=2\times(-4)\times S\\S=\frac{400}{8}=50 m"